OpenStudy (anonymous):

HELP PLEASEE : Solve: x2 – 36 = 0 and x2 = 8x – 12 and show your work. Describe what the solution(s) represent to the graph of each. How are the graphs alike? How are they different?

OpenStudy (anonymous):

Ok, lets see what you have.

OpenStudy (anonymous):

The first equation is what we call a special case quadratic. The reason is that it has the form: \[a^2-b^2=(a+b)(a-b)\] Where a is x and b is 6, i.e: \[x^2-36=(x)^2-(6)^2=(x+6)(x-6)\]

OpenStudy (anonymous):

So if \[(x+6)(x-6) = 0\], the only way for this to be true is if x equals -6 or +6